# Estimating velocities of infectious disease spread through spatio-temporal log-Gaussian Cox point processes

### v1

@@ 3.1.4 Inference
$\Lambda(\textbf{u},t)$ $=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\zeta(\textbf{u},t)\right),$
$\zeta(\textbf{u},t|\theta)$ $\sim\text{Gaussian Process }(\beta,\phi(h_{1},h_{2}|\theta)),$
[…]
where $\phi$ is the covariance function defined in Equations (4) and (5), and $\theta$ is a vector of parameters and hyperparameters.
To fit the model, we use the R-INLA package [Rue et al., 2009] by employing a finite basis function defined on a triangulated mesh of the region of study to approximate the SPDE model in this mesh [Simpson et al., 2015]. This approach introduces a finite-dimensional, continuously specified random field of the form
$Z(\textbf{u},t)=\sum_{i=1}^{n}z_{i}\varphi(\textbf{u},t),$
where $\textbf{z}=(z_{1},\ldots,z_{n})^{t}$ is a multivariate Gaussian random vector and $\{\varphi(\textbf{u},t)\}_{i=1}^{n}$ is a set of linearly independent deterministic basis functions.
@@ 3.2 Velocities
$\nabla\psi(\textbf{u},t)=\begin{pmatrix}\displaystyle\frac{\partial\psi(\textbf{u},t)}{\partial x}\\ \vspace{-0.2cm}\\ \displaystyle\frac{\partial\psi(\textbf{u},t)}{\partial y}\\ \end{pmatrix},\hskip 28.45274pt\mathcal{S}(\textbf{u},t)=\left(\mathcal{S}_{x}(\textbf{u},t),\mathcal{S}_{y}(\textbf{u},t)\right).$
The velocity component can also be decomposed as $\mathcal{S}(\textbf{u},t)=s_{\mathbf{v}}(\textbf{u},t)\mathbf{v}(\textbf{u},t)$ where v is a unit vector that represents the direction of the velocity and $s_{\textbf{v}}$ is its magnitude in the direction v. Under these assumptions Equation (6) can be rewritten as
@@
Now, for a fixed velocity direction v and for a known function $\psi$ we can solve Equation (7) for the magnitude $s_{\textbf{v}}$ that is always positive, then
@@
Notice that the denominator in Equation (8) is the absolute value of the spatial directional derivative of the function $\psi$ in the direction v, i.e. $D_{\textbf{v}}\psi(\textbf{u},t)=\textbf{v}(\textbf{u},v)\cdot\nabla\psi(\textbf{u},t)$. Then, using Cauchy-Schwart inequality, we know that the denominator will be bounded by the norm of the spatial gradient, $\left|\textbf{v}(\textbf{u},v)\cdot\nabla\psi(\textbf{u},t)\right|\leq\lVert\nabla\psi(\textbf{u},t)\rVert$. Therefore, the largest instantaneous change of the function $\psi$ in a neighborhood of $(\textbf{u},t)$ is $\lVert\nabla\psi(\textbf{u},t)\rVert$ and is in the direction $\textbf{v}_{\textbf{max}}=\nabla\psi(\textbf{u},t)/\lVert\nabla\psi(\textbf{u},t)\rVert$. As a result, and because the bounded value is the denominator, the direction $\textbf{v}_{\text{max}}$ also provides the direction of minimum velocity at $(\textbf{u},t)$ that now could be written as
@@
3.2.1 Computing velocities
We estimate the minimal velocity $s_{\text{min}}(\textbf{u},t)$ using the output of the model described in Section 3.1.4 which provides the estimated intensity function $\hat{\lambda}(\textbf{u},t)$ over a spatio-temporal grid.
@@ 3.2.1 Computing velocities
Let $\Delta_{t}$, $\Delta_{x}$, $\Delta_{y}$ denote the discrete step size in time $t$, latitude $x$ and longitude $y$ respectively, and let $\hat{\lambda}(i,j,n)$ be the estimated intensity function in the grid indexed in space for $(i,j)$ and time $n$. Then, the derivative of the intensity in time that appears in the numerator of $s_{\text{min}}(\textbf{u},t)$ can be approximated using numerical methods, such as the symmetric difference quotient
$\frac{\partial\hat{\lambda}(\textbf{u},t)}{\partial t}\approx\frac{\hat{\lambda}(i,j,n+1)-\hat{\lambda}(i,j,n-1)}{2\Delta_{t}}.$ (11)
@@
Notice that the temporal derivative approximation presented in Equation (11) needs future values in time. However, if we are interested in predictions, we can use a first-order divided difference if we do not have access to this value.
@@
To estimate the denominator $\lVert\nabla\lambda(\textbf{u},t)\rVert$, following [Sethian, 1999], we first define the forward and backward difference operators in space as follows
@@
An approximation of the norm of the spatial gradient is then obtained as the average the last four expressions
@@
Finally, the minimal velocity described in Equation (10) can be approximated using the approximations described in Equations (12) and (11) as follows
$s_{\text{min}}(\textbf{u},t)\approx\frac{\left|\displaystyle\frac{\hat{\lambda}(i,j,n+1)-\hat{\lambda}(i,j,n-1)}{2\Delta_{t}}\right|}{\left(\frac{1}{4}\right)\displaystyle\sum_{k=1}^{4}\lVert\nabla_{k}\hat{\lambda}(i,j,n)\rVert}.$ (13)
4 Simulations
We perform a simulation study to assess the performance of our model in comparison with an underlying truth intensity function. Specifically, we consider a deterministic intensity function to generate a spatio-temporal point pattern in the unit cube, $(\textbf{u},t)\in[0,1]^{3}$, defined by
@@ 4 Simulations
This intensity function, proposed by [Schliep and Gelfand, 2019], was chosen because its close form allows us to compute partial derivatives in space and time, and with that, we can compute the true minimal velocity described in Equation (10).
The partial derivatives of this intensity function has the following form
$\frac{\partial\lambda(x,y,t)}{\partial x}$ $=\lambda(x,y,t)\left[(1-t)\beta_{1}f_{1x}(x,y)+t(1-t)\beta_{2}f_{2x}(x,y)+t^{2}\beta_{3}f_{3x}\right],$
$\frac{\partial\lambda(x,y,t)}{\partial y}$ $=\lambda(x,y,t)\left[(1-t)\beta_{1}f_{1y}(x,y)+t(1-t)\beta_{2}f_{2y}(x,y)+t^{2}\beta_{3}f_{3y}\right],$
$\frac{\partial\lambda(x,y,t)}{\partial t}$ $=\lambda(x,y,t)\left[-\beta_{1}f_{1}(x,y)+(1-2t)\beta_{2}f_{2}(x,y)+2t\beta_{3}f_{3}\right],$
where for $k=1,2,3$, $f_{kx}$ denotes the partial derivative in the $x$ direction of $f_{k}$ and have the following closed form
$f_{kx}(x,y)$ $=\frac{\partial f_{k}(x,y)}{\partial x}$
$=f_{k}(x,y)\left(-(x-\mu_{k1})\Sigma_{k,11}^{-1}-(y-\mu_{k2})\Sigma_{k,12}^{-1}\right),$
and $\Sigma_{k,ij}^{-1}$ is the $i$th row and $j$th column of $\Sigma_{k}^{-1}$. Analogously, the partial derivative of the density function $f_{k}$ in the direction $y$ is
$f_{ky}(x,y)$ $=\frac{\partial f_{k}(x,y)}{\partial y}$
$=f_{k}(x,y)\left(-(x-\mu_{k1})\Sigma_{k,12}^{-1}-(y-\mu_{k2})\Sigma_{k,22}^{-1}\right).$
To simulate the spatio-temporal point pattern, we fixed the parameters as $\lambda_{0}=100$, $\beta_{0}=-1.5$, $\beta_{1}=8.0$, $\beta_{2}=2.0$, $\beta_{3}=2.0$, $\mu_{1}=(0.4,0.2)$, $\mu_{2}=(0.8,0.5)$, $\mu_{3}=(0.2,0.8)$, and
@@
With these fixed parameters, we use the R package spatstat [Baddeley et al., 2015] to generate a realization of an inhomogeneous Poisson process with intensity function given by Equation (14) that generates 7840 points distributed in a grid of temporal resolution of $1/20$ and spatial resolution of $1/30$. These parameters were chosen to show spatio-temporal variability in the intensity function such that the process starts with a mode in $\mu_{1}$, for $t\in(0.0,5)$ the mode will move from $\mu_{1}$ to $\mu_{2}$, and for $t\in(0,1)$ the mode will move from $\mu_{2}$ to $\mu_{3}$ to finally finish with a mode in $\mu_{3}$. In Figure 1, we can see how the mode of the intensity function varies in time points $0,0.25,0.5,0.75$, and $1$.
Figure 1: Variation of the intensity in times $0,0.25,0.5,0.75,$ and $1$.
Figure 2 compares the true intensity function and the posterior mean generated by the model at time points $0.225$, $0.575$, and $0.875$. As expected, the model captures the movement of the intensity mode across time. Moreover, we can see that the posterior mean differs slightly from the true values near the boundaries. This happens mainly because the boundaries were artificially defined in the unit cube. However, in general, we can see that the model effectively captures the overall changes in intensity across time.
True and computed spatial and temporal derivatives at time $t=0.225$ are shown in Figure 3. For computing temporal derivatives in times $0.225$ and $0.575$, we use a step size of $\Delta_{t}=0.2$. When $t=0.875$, we use a smaller step size of $\Delta_{t}=0.1$ in time because the spatio-temporal point process is defined in time in the interval $[0,1]$. Similarly, true and approximation of the intensity partial derivatives for times $0.575$ and $0.875$ are shown in Figures 4 and 5. Overall, we can see that the approximation using finite differences captures the trend, maximum, and minimum values of the true values.
(a) $t=0.225$
(b) $t=0.575$
(c) $t=0.875$
Figure 2: True intensity function (left) and posterior mean (right) for the time points $t=0.225$, $t=0.575$ and $t=0.875$.
(a) $\frac{\partial\lambda(\textbf{u},t)}{\partial x}$
(b) $\frac{\partial\lambda(\textbf{u},t)}{\partial y}$
(c) $\frac{\partial\lambda(\textbf{u},t)}{\partial t}$
Figure 3: True partial derivatives of the intensity function (left) and finite differences estimation of the partial derivatives (right) at $t=0.225$
(a) $\frac{\partial\lambda(\textbf{u},t)}{\partial x}$
(b) $\frac{\partial\lambda(\textbf{u},t)}{\partial y}$
(c) $\frac{\partial\lambda(\textbf{u},t)}{\partial t}$
Figure 4: True partial derivatives of the intensity function (left) and finite differences estimation of the partial derivatives (right) at $t=0.575$
(a) $\frac{\partial\lambda(\textbf{u},t)}{\partial x}$
(b) $\frac{\partial\lambda(\textbf{u},t)}{\partial y}$
(c) $\frac{\partial\lambda(\textbf{u},t)}{\partial t}$
Figure 5: True partial derivatives of the intensity function (left) and finite differences estimation of the partial derivatives (right) at $t=0.875$
Using the partial derivatives computed before and the gradient approximation described in Equation (12), we can approximate the minimum velocity of the spatio-temporal point process using Equation (13). The true minimum velocity and its approximation for times $0.225$, $0.575$, and $0.875$ are shown in Figure 6. In this figure, the arrows are the standardized vector that shows the direction in which the slowest speed of change in the intensity is obtained and its positive, i.e.,
$\vec{s}=\text{sign}\left(\frac{\partial\lambda(\textbf{u},t)}{\partial t}\right)\cdot\frac{\nabla\psi(\textbf{u},t)}{\lVert\nabla\psi(\textbf{u},t)\rVert},$ (15)
where the sign is a function that is $-1$ if the number is negative, 00 if it is zero, or $+1$ if it is positive.
(a) $t=0.225$
(b) $t=0.575$
(c) $t=0.875$
Figure 6: True minimum velocity of the intensity function (left) and finite difference approximation of this for the time points $0.225$, $0.575$, and $0.875$. The arrows presented show the slowest speed of change and the direction of this speed is obtained and is positive.
5 Application

### v2

@@ 3.1.4 Inference
$\begin{split}\Lambda(\textbf{u},t)&=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\xi(\textbf{u},t)\right),\\ \xi(\textbf{u},t|\theta,\alpha)&\sim\text{Gaussian process }(\beta,\phi(h_{1},h_{2}|\theta,\alpha)),\\ \alpha=(\alpha_{t},\alpha_{u},\alpha_{e})&\sim\text{Smoothing parameters},\\ \theta&\sim\text{priors}\end{split}$ (6)
where $\phi$ is the covariance function that comes with the solution of the spatio-temporal SPDE defined by (5).
To fit the model, we use the INLAspacetime package [lindgren2023inlaspacetime] by employing a finite basis function defined on a triangulated mesh of the region of study to approximate the spatio-temporal SPDE model described in (5) in this mesh [simpson_et_al_2015].
As discussed in Section 3.1.3, the smoothing parameters regulate the smoothness of the Gaussian process in both space and time, as well as its type of separability. In summary, Table 1 presents the spatio-temporal models available in the INLAspacetime package, their type of separability, and the corresponding temporal and spatial smoothness parameters.
Model $\alpha_{t}$ $\alpha_{s}$ $\alpha_{e}$ Type $v_{t}$ $v_{s}$
A 1 0 2 Separable order 1 $1/2$ 1
B 1 2 1 Critical diffusion $1/2$ 1
C 2 0 2 Separable order 2 $3/2$ 1
D 2 2 0 Iterated diffusion 1 2
Table 1: Models available in the R-package INLAspacetime.
In Table 1, Model B (Critical diffusion) is nonseparable with a separability coefficient of $\beta_{u}=0.5$. At the same time, Model D (Iterated diffusion) is fully non-separable with $\beta_{u}=0$. Since velocity estimation requires computing first derivatives in both space and time, the intensity function of the LGCP process must be at least once differentiable. We focus on Models C and D to satisfy this requirement, ensuring the necessary smoothness. In Section 4, we evaluate and compare their performance.
@@ 3.2 Velocities
$\nabla\psi(\textbf{u},t)=\begin{pmatrix}\displaystyle\frac{\partial\psi(\textbf{u},t)}{\partial x}\\ \vskip-5.69046pt\\ \displaystyle\frac{\partial\psi(\textbf{u},t)}{\partial y}\\ \end{pmatrix},\hskip 28.45274pt\mathcal{S}(\textbf{u},t)=\left(\mathcal{S}_{x}(\textbf{u},t),\mathcal{S}_{y}(\textbf{u},t)\right).$
The velocity component can also be decomposed as $\mathcal{S}(\textbf{u},t)=s_{\mathbf{v}}(\textbf{u},t)\mathbf{v}(\textbf{u},t)$ where v is a unit vector that represents the direction of the velocity and $s_{\textbf{v}}$ is its magnitude in the direction v. Under these assumptions, (7) can be rewritten as
@@
Now, for a fixed velocity direction v and for a known function $\psi$, we can solve (8) for the magnitude $s_{\textbf{v}}$ that is always positive. Then
@@
Notice that the denominator in (9) is the absolute value of the spatial directional derivative of the function $\psi$ in the direction v, i.e. $D_{\textbf{v}}\psi(\textbf{u},t)=\textbf{v}(\textbf{u},v)\cdot\nabla\psi(\textbf{u},t)$. Then, using Cauchy-Schwart inequality, we know that the denominator will be bounded by the norm of the spatial gradient, $\left|\textbf{v}(\textbf{u},v)\cdot\nabla\psi(\textbf{u},t)\right|\leq\lVert\nabla\psi(\textbf{u},t)\rVert$. Therefore, the largest instantaneous change of the function $\psi$ in a neighborhood of $(\textbf{u},t)$ is $\lVert\nabla\psi(\textbf{u},t)\rVert$ and is in the direction $\textbf{v}_{\textbf{max}}=\nabla\psi(\textbf{u},t)/\lVert\nabla\psi(\textbf{u},t)\rVert$. As a result, and because the bounded value is the denominator, the direction $\textbf{v}_{\text{max}}$ also provides the direction of minimum velocity at $(\textbf{u},t)$ that now could be written as
@@
The direction of this minimal velocity i.e., the direction of the slowest rate of intensity change, which remains positive i.e.d.
$\vec{s}_{\text{min}}(\textbf{u},t)=\text{sign}\left(\frac{\partial\lambda(\textbf{u},t)}{\partial t}\right)\cdot\frac{\nabla\lambda(\textbf{u},t)}{\lVert\nabla\lambda(\textbf{u},t)\rVert}.$ (12)
where the sign is a function that is 1-1 if the number is negative, 0 if it is zero, or +1+1 if it is positive. This vector captures both the magnitude and direction of the slowest rate of change of intensity, making it a useful quantity for interpreting in the context of disease spread, how and where events are emerging or disappearing in space-time.
3.3 Using velocities to understand disease spread
In this section, we present illustrative examples of intensity functions that evolve over time, along with their corresponding velocity fields. These examples help clarify how to interpret both the magnitude and direction of velocity in the context of infectious disease dynamics. Specifically, we consider three scenarios: one in which a hotspot emerges, another in which it disperses, and a third in which the hotspot shifts its location horizontally. All of the examples are defined spatially in the unit square $0,1]^{2}$ and temporally in the unit interval $[0,1]$, and all the velocities are computed at the middle point at $t=0.5$. Also, the examples have the LGCP structure, so it is the exponential of a linear combination of the density of a bivariate normal distribution with a diagonal covariance matrix.
(a) Example 1
(b) Example 2
(c) Example 3
Figure 1: Magnitude and direction of the true minimal velocities at $t=0.5$. Subfigure (a) illustrates the emergence of a hotspot, (b) shows the disappearance of a hotspot, and (c) depicts the horizontal movement of a hotspot.
In the first scenario, the intensity is zero at time $t=0$ and gradually increases over time. For each $t>0$, the intensity has a single mode centered at $(0.5,0.5)$, reaching its maximum at $t=1$. This setup models the emergence of a central hotspot. As shown in Figure 1 (a), at $t=0.5$, the velocity is strongest in the center of the domain, with arrows pointing inward toward the hotspot. This inward flow produces a plughole-like vortex effect, indicating the formation and growth of a localized outbreak.
In the second scenario, the intensity pattern is reversed: it starts with a central hotspot at $(0.5,0.5)$ at $t=0$ with maximum intensity, which gradually decreases until it vanishes at $t=1$. This represents a scenario where an initially concentrated outbreak becomes controlled and dissipates over time. As seen in Figure 1 (b), the velocity magnitude is again highest in the central region; however, unlike the first case, the arrows point outward from the hotspot, producing a seed-dispersal-like effect that reflects the spatial diffusion or dissolution of the outbreak.
In the third scenario, a hotspot initially located at $(0.2,0.5)$ at $t=0$ shifts horizontally to $(0.8,0.5)$ by $t=1$. The mode remains strong throughout, but its spatial location changes. This setup models a situation where the outbreak relocates from one part of the region to another, indicating spatial movement of disease activity. In Figure 1 (c), the velocity field shows two prominent zones: one on the left and one on the right. The left-hand hotspot displays outward-pointing arrows, signaling the disintegration of the hotspot. In contrast, the right-hand hotspot has inward-pointing arrows, indicating the accumulation of cases and emergence of a new hotspot.
Notice that in more complex scenarios, when there are more than two hotspots, there could also be scenarios of dispersion in zones of high velocity magnitude and arrows pointing to the same direction, as shown in the simulation study for $t=0.575$ in Figure 3 in the second row, first column.
3.4 Computing velocities
@@
We estimate the minimal velocity $s_{\text{min}}(\textbf{u},t)$ using the output of the model described in Section 3.1.4 which provides the estimated intensity function $\hat{\lambda}(\textbf{u},t)$ over a spatio-temporal grid. Let $\Delta_{t}$, $\Delta_{x}$, $\Delta_{y}$ denote the discrete step size in time $t$, latitude $x$ and longitude $y$ respectively, and let $\hat{\lambda}(i,j,n)$ be the estimated intensity function in the grid indexed in space for $(i,j)$ and time $n$. Then, the derivative of the intensity in time that appears in the numerator of $s_{\text{min}}(\textbf{u},t)$ can be approximated using numerical methods, such as backward differentiation
$\frac{\partial\hat{\lambda}(\textbf{u},t)}{\partial t}\approx\frac{\hat{\lambda}(i,j,n)-\hat{\lambda}(i,j,n-1)}{\Delta_{t}}.$ (13)
@@
Notice that the temporal derivative approximation presented in (13) needs future values in time. However, if we are interested in predictions, we can use a first-order divided difference if we do not have access to this value.
@@
To estimate the denominator $\lVert\nabla\lambda(\textbf{u},t)\rVert$, following [sethian1999], we first define the forward and backward difference operators in space as follows
@@
An approximation of the norm of the spatial gradient is then obtained as the average of the last four expressions
@@
Finally, the minimal velocity described in Equation (11) can be approximated using the approximations described in (14) and (13), as follows
$s_{\text{min}}(\textbf{u},t)\approx\frac{\left|\displaystyle\frac{\hat{\lambda}(i,j,n)-\hat{\lambda}(i,j,n-1)}{\Delta_{t}}\right|}{\left(\frac{1}{4}\right)\displaystyle\sum_{k=1}^{4}\lVert\nabla_{k}\hat{\lambda}(i,j,n)\rVert}.$ (15)
4 Simulation study
In this section, we perform a simulation study to assess the performance of our model in comparison with an underlying truth intensity function. Specifically, we consider a deterministic intensity function to generate a spatio-temporal point pattern in the unit cube, $(\textbf{u},t)\in[0,1]^{3}$, defined by
@@ 4 Simulation study
This intensity function, proposed in [schliep_gelfand_2019], was chosen because its close form allows to compute partial derivatives in space and time, and with that, we can compute the true minimal velocity described in (11).
The partial derivatives of this intensity function have the following form
$\begin{split}\frac{\partial\lambda(x,y,t)}{\partial x}&=\lambda(x,y,t)\left[(1-t)\beta_{1}f_{1x}(x,y)+t(1-t)\beta_{2}f_{2x}(x,y)+t^{2}\beta_{3}f_{3x}\right],\\ \frac{\partial\lambda(x,y,t)}{\partial y}&=\lambda(x,y,t)\left[(1-t)\beta_{1}f_{1y}(x,y)+t(1-t)\beta_{2}f_{2y}(x,y)+t^{2}\beta_{3}f_{3y}\right],\\ \frac{\partial\lambda(x,y,t)}{\partial t}&=\lambda(x,y,t)\left[-\beta_{1}f_{1}(x,y)+(1-2t)\beta_{2}f_{2}(x,y)+2t\beta_{3}f_{3}\right],\end{split}$ (17)
where for $k=1,2,3$, $f_{kx}$ denotes the partial derivative in the $x$ direction of $f_{k}$, and have the following closed form
$\begin{split}f_{kx}(x,y)&=\frac{\partial f_{k}(x,y)}{\partial x}\\ &=f_{k}(x,y)\left(-(x-\mu_{k1})\Sigma_{k,11}^{-1}-(y-\mu_{k2})\Sigma_{k,12}^{-1}\right),\end{split}$ (18)
and $\Sigma_{k,ij}^{-1}$ is the $i$th row and $j$th column of $\Sigma_{k}^{-1}$. Analogously, the partial derivative of the density function $f_{k}$ in direction $y$ is
$\begin{split}f_{ky}(x,y)&=\frac{\partial f_{k}(x,y)}{\partial y}\\ &=f_{k}(x,y)\left(-(x-\mu_{k1})\Sigma_{k,12}^{-1}-(y-\mu_{k2})\Sigma_{k,22}^{-1}\right).\end{split}$ (19)
To simulate the spatio-temporal point pattern, we fixed the parameters as $\beta_{0}=-1.5$, $\beta_{1}=8.0$, $\beta_{2}=2.0$, $\beta_{3}=2.0$, $\mu_{1}=(0.4,0.2)$, $\mu_{2}=(0.8,0.5)$, $\mu_{3}=(0.2,0.8)$, and
@@
Once we fixed the other parameters, for $\lambda_{0}$, which controls the number of points in the simulation, we proposed the following scenarios: $\lambda_{0}\in[5,10,20,30]$.
@@
With these fixed parameters, we use the R package spatstat [baddeley_2015] to generate realizations of an inhomogeneous Poisson process with intensity function given by (16) for each $\lambda_{0}$. This results in 359, 773, 1526, and 2276 points distributed on a grid with a temporal resolution of $1/20$ and a spatial resolution of $1/40$. These settings were chosen to capture spatio-temporal variability in the intensity function. Specifically, the process begins with a mode at $\mu_{1}$, then shifts from $\mu_{1}$ to $\mu_{2}$ for $t\in(0,5)$, moves from $\mu_{2}$ to $\mu_{3}$ for $t\in(0,1)$, and finally stabilizes at $\mu_{3}$.
Figure 2 illustrates the intensity function for $\lambda_{0}=10$, showing how the mode evolves over time at points $0,0.25,0.5,0.75$, and $1$. Note that $\lambda_{0}$ affects only the magnitude of the intensity function, leaving the mode unchanged across all cases.
Time = 0
Time = 0.25
Time = 0.5
Time = 0.75
Time = 1
Figure 2: Temporal evolution of the simulated intensity function at times $t=0$, $0.25$, $0.5$, $0.75$, and $1$. Each panel shows the spatial distribution and variation of intensity over time, illustrating the movement of the mode over time.
To illustrate the computation of velocity, we used three time points: $0.225$, $0.575$, and $0.875$. These were selected to capture key transition phases. First, $0.225$ marks the beginning of the change from the first mode $\mu_{1}$ to the second $\mu_{2}$, $0.575$ corresponds with the "middle" point of connections between the three modes, and $0.875$ represents the final shift from $\mu_{2}$ to $\mu_{3}$.
Based on these times, Table 2 presents the root mean squared error (RMSE) between the velocity estimates obtained using (15) and the true values from (17), (18), and (19) at times $0.225$, $0.575$, and $0.875$. The minimum RMSE for each case is highlighted in gray, comparing the LGCP model with a separable covariance structure (Model C) and the fully nonseparable model (Model D). The nonseparable model usually achieves a better fit, yielding a lower RMSE than its separable counterpart. This trend is particularly evident when the number of points is low ($\lambda_{0}=5$), where the fully nonseparable model is better than the separable model across all time points considered. This characteristic is especially relevant for epidemiological data, where case numbers are typically low in the early stages of disease spread. In such scenarios, the improved accuracy of the nonseparable model makes it a valuable tool for better capturing and analyzing disease dynamics.
Time $\lambda_{0}$ Model C Model D
0.225 5 0.6959 0.5154
0.575 5 0.5866 0.5302
0.875 5 1.8021 0.6843
0.225 10 0.6216 0.4917
0.575 10 0.5098 0.5511
0.875 10 0.8678 0.8721
0.225 20 0.6366 0.5754
0.575 20 0.5006 0.5654
0.875 20 0.9694 0.9552
0.225 30 0.8499 0.5480
0.575 30 0.6003 0.5273
0.875 30 0.7827 0.8578
Table 2: Root mean squared error (RMSE) of the estimated velocities for the separable model (Model C) and the non-separable model (Model D), compared to the true values. For each configuration, the lower RMSE between the two models is highlighted in grey.
In Appendix A, we present the estimation of the intensity function using INLAspacetime and inlabru for time points $0.225$, $0.575$, and $0.875$, with $\lambda_{0}=5$. We also describe the approximation of the absolute value of the temporal derivative and the norm of the spatial gradient of the intensity function, computed using finite differences as outlined in Section 3.4. The temporal derivative is approximated using backward differences with a step size of $\Delta_{t}=0.2$, while the spatial partial derivatives are calculated using both forward and backward differences with a step size of $\Delta_{x}=\Delta_{y}=1/40$.
Velocity of the spatio-temporal point pattern
True Values
Model C
Model D
Time = 0.225
Time = 0.575
Time = 0.875
Figure 3: Velocity estimation for the simulated spatio-temporal point pattern. Columns correspond to the true velocity (left), the estimate obtained using the separable model (Model C, center), and the estimate from the nonseparable model (Model D, right). Rows represent three key time points in the process: 0.225 (top), 0.575 (middle), and 0.875 (bottom). Each panel shows the magnitude and direction of the minimal velocity.
Figure 3 presents the true and estimated velocities at time points $0.225$, $0.575$, and $0.875$ for $\lambda_{0}=5$. The velocity estimates are obtained from the posterior mean of the diffusion model, where partial derivatives in time and space are computed and subsequently used in (15) to derive minimal velocities. The results indicate that the non-separable model (Model D) generally provides a more accurate velocity estimation compared to the Separable model (Model C) at time points $0.225$ and $0.875$. Additionally, the non-separable (Model D) yields a smoother velocity field across all time points, particularly at $t=0.575$, where it significantly reduces noise compared to the separable model (Model C). Finally, the arrows in Figure 3 are computed according to equation (12), where the sign of the temporal derivative determines orientation, and the gradient of the intensity defines the direction of steepest spatial change.
All code used for this simulation study is publicly available at https://github.com/fravellaneda/estimating-velocities.
5 COVID-19 spread and velocities in Colombia